how do I graph f(x)=-3x2 + 8x
Answer by nyc_function(2741)

This is an upside down parabola.
Just select about 6 different values for x and then simplify to find y values that together make points in the form (x,y).
Let y = f(x).
y = -3x^2 + 8x
For example, if x = 0, find y.
y = -3(0)^2 + 8(0)
y = 0 + 0
y = 0
This makes the point (0,0). You can now plot this point on the xy-plane.
Find other points in the same way. When you have about 5 or 6 points, plot each point and then connect them with a smooth curve.
Can you finish now?